OpenStudy (anonymous):

Divide. x^2-4/x-8/x-2

OpenStudy (anonymous):

\[\frac{\frac{x^2-4}{x-8}}{x-2}\]?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

same as \[\frac{x^2-4}{(x-8)(x-2)}\] factor the numerator, and cancel the common factor of \(x-2\) top and bottom

OpenStudy (anonymous):

as in \[\frac{(x+2)(x-2)}{(x-8)(x-2)}\] then cancel

OpenStudy (anonymous):

so x+2/x-8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok ty satellite73